# The entries of the Sinkhorn limit of an $m\times n$ matrix

### v1

@@ Abstract.
We use Gröbner bases to compute the Sinkhorn limit of a positive $3\times 3$ matrix $A$, showing that the entries are algebraic numbers with degree at most $6$. The polynomial equation satisfied by each entry is large, but we show that it has a natural representation in terms of linear combinations of products of minors of $A$. We then use this representation to interpolate a polynomial equation satisfied by an entry of the Sinkhorn limit of a positive $4\times 4$ matrix. Finally, we use the PSLQ algorithm and $1.5$ years of CPU time to formulate a conjecture, up to certain signs that we have not been able to identify, for a polynomial equation satisfied by an entry of the Sinkhorn limit of a positive $m\times n$ matrix. In particular, we conjecture that the entries are algebraic numbers with degree at most $\binom{m+n-2}{m-1}$. This degree has a combinatorial interpretation as the number of minors of an $(m-1)\times(n-1)$ matrix, and the coefficients reflect new combinatorial structure on sets of minor specifications.
@@
In a 1964 paper, Sinkhorn [19] considered the following iterative scaling process. Let $A$ be a square matrix with positive entries. Scale the rows so that each row sum is $1$. Then scale the columns so that each column sum is $1$; generically, this changes the row sums. To restore the row sums $1$, scale the rows again, then scale the columns, and so on. Sinkhorn showed that the sequence of matrices obtained through this process converges to a matrix whose row and column sums are $1$ (in other words, a doubly stochastic matrix). We call this matrix the Sinkhorn limit of $A$ and denote it $\operatorname{Sink}(A)$. Sinkhorn also showed that $\operatorname{Sink}(A)$ is the unique doubly stochastic matrix $S$ with the same size as $A$ such that $S=RAC$ for some diagonal matrices $R$ and $C$ with positive diagonal entries. Here $R$ can be taken to be the product of the row-scaling matrices and $C$ the product of the column-scaling matrices.
The literature on questions related to the iterative scaling process is large; Idel’s extensive survey [10] covers results as of 2016. Mathematical applications include preconditioning linear systems to improve numerical stability, approximating the permanent of a matrix, and determining whether a graph has a perfect matching. In a number of other areas, iterative scaling was discovered independently multiple times [14, 6, 2], and it is used in machine learning to efficiently compute optimal transport distances [5]. As a result of its ubiquity and importance, many authors have been interested in fast algorithms for approximating Sinkhorn limits numerically [9, 11, 12, 15, 13, 1, 4].
@@
In particular, the top left entry $x$ of $\operatorname{Sink}(A)$ satisfies
@@
For general $3\times 3$ matrices, Chen and Varghese [3] used numeric experiments to conjecture that the entries of $\operatorname{Sink}(A)$ generically have degree $6$ over the field generated by the entries of $A$.
@@
Let $x$ be the top left entry of $\operatorname{Sink}(A)$. The PSLQ integer relation algorithm [8] can be used to recognize an algebraic number, given a sufficiently high-precision approximation. For the approximation
@@
In Section 3, we use the form of the equation for $3\times 3$ matrices to infer the form of the equation for general $n\times n$ matrices $A$. In particular, we conjecture that the entries of $\operatorname{Sink}(A)$ generically have degree $\binom{2n-2}{n-1}$. For a general $4\times 4$ matrix, the Gröbner basis computation is infeasible, so instead we use PSLQ to recognize entries of Sinkhorn limits for enough integer matrices to solve for each coefficient in the equation. We also obtain several coefficients in the equation for $5\times 5$ matrices with this method.
@@
In Section 4, we generalize to non-square matrices by defining the Sinkhorn limit of an $m\times n$ matrix $A$ to be the matrix obtained by iteratively scaling so that each row sum is $1$ and each column sum is $\frac{m}{n}$. Again we use PSLQ and solve for coefficients in equations for matrices of various sizes. We then interpolate formulas for these coefficients as functions of $m$ and $n$. By identifying combinatorial structure in these formulas, we build up to the following conjecture. Let $D(m,n)$ be the set of all pairs $(R,C)$ where $R\subseteq\{2,3,\dots,m\}$, $C\subseteq\{2,3,\dots,n\}$, and $\lvert{R}\rvert=\lvert{C}\rvert$. The set $D(m,n)$ consists of the specifications of all minors of an $m\times n$ matrix $A$ that do not involve the first row or first column. For each $S\subseteq D(m,n)$, let $M(S)$ be the product of minors of $A$ defined in Section 2. Let $\operatorname{adj}_{S,\sigma(S)}(m,n)$ be the $\lvert{S}\rvert\times\lvert{S}\rvert$ matrix defined in Section 4; this matrix resembles an adjacency matrix, and its entries are linear functions of $m$ and $n$ with signs determined by the sign alteration $\sigma(S)$.
@@
Let $m\geq 1$ and $n\geq 1$. For each $S\subseteq D(m,n)$, there exists a sign alteration $\sigma(S)$ such that, for every $m\times n$ matrix $A$ with positive entries, the top left entry $x$ of $\operatorname{Sink}(A)$ satisfies
$\sum_{S\subseteq D(m,n)}\left(\det\operatorname{adj}_{S,\sigma(S)}(m,n)\right)M(S)x^{\lvert{S}\rvert}=0.$
@@
We conclude in Section 5 with several open questions. In particular, we mention that a more general process of iteratively scaling was introduced in 1937 by Kruithof in the context of predicting telephone traffic [14, Appendix 3d]. We conjecture that the entries of this more general limit also have degree at most $\binom{m+n-2}{m-1}$.
Our Mathematica package SinkhornPolynomials [18] uses the results of this paper to compute Sinkhorn limits rigorously for $3\times 3$ matrices and conjecturally for $2\times n$, $3\times 4$, $3\times 5$, and $4\times 4$ matrices as well as their transposes.
@@ 2. The Sinkhorn limit of a $3\times 3$ matrix
It suffices to describe the top left entry of $\operatorname{Sink}(A)$. This is because the iterative scaling process isn’t sensitive to the order of the rows or the order of the columns. Therefore, if $R$ is the permutation matrix swapping rows $1$ and $i$ and $C$ is the permutation matrix swapping columns $1$ and $j$, then $R\operatorname{Sink}(A)C=\operatorname{Sink}(RAC)$. In particular, the $(i,j)$ entry of $\operatorname{Sink}(A)$ is equal to the $(1,1)$ entry of $\operatorname{Sink}(RAC)$. For example, the $(2,3)$ entry of $\operatorname{Sink}(A)$ is equal to the $(1,1)$ entry of
$\operatorname{Sink}\!\left(\begin{bmatrix}a_{23}&a_{22}&a_{21}\\ a_{13}&a_{12}&a_{11}\\ a_{33}&a_{32}&a_{31}\end{bmatrix}\right).$
To compute a polynomial equation satisfied by the top left entry of $\operatorname{Sink}(A)$, we set up three matrices
@@
Let $A$ be a positive $3\times 3$ matrix. The top left entry $x$ of $\operatorname{Sink}(A)$ satisfies $b_{6}x^{6}+\dots+b_{1}x+b_{0}=0$, where the coefficients $b_{k}$ appear in Table 1 in factored form.
@@
Table 1. Coefficients in the polynomial equation satisfied by the top left entry of the Sinkhorn limit of a positive $3\times 3$ matrix.
In particular, the degree of each entry of $\operatorname{Sink}(A)$ is at most $6$. For the matrix $A$ in Example 1, Theorem 3 states that the top left entry $x$ of $\operatorname{Sink}(A)$ satisfies
@@
This agrees with Equation (2). This polynomial is irreducible, and this confirms the conjecture of Chen and Varghese [3] that the entries of $\operatorname{Sink}(A)$ for positive $3\times 3$ matrices $A$ generically have degree $6$.
Let $f(x)=b_{6}x^{6}+b_{5}x^{5}+\dots+b_{1}x+b_{0}$ be the polynomial in Theorem 3. Project $A$ to a symmetric matrix by setting $a_{21}=a_{12}$, $a_{31}=a_{13}$, and $a_{32}=a_{23}$. Then the projection of $f(x)$ factors as $-\left((a_{11}a_{23}-a_{12}a_{13})x-a_{11}a_{23}\right)^{2}g(x)$, where $g(x)$ is the degree-$4$ polynomial computed by Ekhad and Zeilberger [7] for the top left entry.
An obvious question is whether there is a better way to write the polynomial $f(x)$ in Theorem 3. In fact there is, using determinants. We first observe that $b_{k}$ contains the factor $a_{11}^{5-k}$ for each $k\in\{0,1,\dots,5\}$. This suggests that the scaled polynomial $a_{11}f(x)$ is more natural than $f(x)$, since the coefficient of $x^{k}$ in $a_{11}f(x)$ contains the factor $a_{11}^{6-k}$ not just for $k\in\{0,1,\dots,5\}$ but for all $k\in\{0,1,\dots,6\}$ (where in fact the coefficient of $x^{6}$ also contains $a_{11}^{1}$).
Now the leading coefficient $a_{11}b_{6}$ of $a_{11}f(x)$ is the product of $6$ minors of $A$. More specifically, it is the product of all the minors of $A$ that involve $a_{11}$. Furthermore, the constant coefficient $a_{11}b_{0}$ of $a_{11}f(x)$ is the product of $a_{11}^{6}$ and the $6$ minors of $A$ that do not involve the first row or first column (including the determinant $1$ of the $0\times 0$ matrix).
@@
We have seen that this is the case for $a_{11}b_{6}$ and $a_{11}b_{0}$. Remarkably, it turns out that $a_{11}b_{5},a_{11}b_{4},\dots,a_{11}b_{1}$ can be written this way as well. Some notation for these minors will be useful.
Notation.
Define
Define $A_{R,C}$ to be the submatrix of $A$ obtained by extracting the rows indexed by $R$ and the columns indexed by $C$. For each $(R,C)\in D(m,n)$, define
The minor $\Delta\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)$ involves $a_{11}$, and $\Gamma\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)$ is the product of $a_{11}$ and a minor that does not involve the first row or first column. This notation does not reflect the dependence on $A$, but the matrix will be clear from context. Each subset $S\subseteq D(m,n)$ specifies a monomial in the expressions $\Delta\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)$ and $\Gamma\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)$, namely
Each element of $D(m,n)$ contributes to the monomial $M(S)$ as the argument of either $\Delta$ or $\Gamma$; indeed we specify each monomial by the elements that appear as arguments of $\Delta$. To make a subset $S=\{(R_{1},C_{1}),(R_{2},C_{2}),\dots,(R_{k},C_{k})\}$ easier to read, we format it as
@@
We continue to let $A$ be the general $3\times 3$ matrix in Equation (3), and let $x$ be the top left entry of $\operatorname{Sink}(A)$. To rewrite the coefficients $a_{11}b_{k}$, we look for coefficients $c_{S}\in\mathbb{Q}$, indexed by subsets $S\subseteq D(3,3)$, such that
@@
Therefore the expression for $b_{3}$ in Table 1 does not uniquely determine the coefficients $c_{S}$ where $\lvert{S}\rvert=3$. However, there is additional information we can use to obtain uniqueness. The polynomial $f(x)$ possesses symmetries arising from the following invariance properties.
Since the iterative scaling process isn’t sensitive to row order, the top left entry of $\operatorname{Sink}(A)$ is invariant under row permutations of $A$ that fix the first row. Similarly for column permutations. Additionally, the top left entry of $\operatorname{Sink}(A)$ is invariant under transposition of $A$; this follows from Sinkhorn’s result that $\operatorname{Sink}(A)$ is the unique doubly stochastic matrix $S$ such that $S=RAC$ for some diagonal matrices $R$ and $C$. This suggests the following equivalence relation.
@@
Motivated by Example 5, we add the constraint that $c_{T}=c_{S}$ when $T\equiv S$, so that the coefficients $c_{S}$ reflect the symmetries of $f(x)$. With this constraint, the coefficient $a_{11}b_{3}$ has a unique representation as a linear combination of monomials $M(S)$ where $\lvert{S}\rvert=3$. Writing each coefficient in Table 1 using class sums gives the following improvement of Theorem 3. Here $d_{k}=a_{11}b_{k}$.
@@
Let $A$ be a positive $3\times 3$ matrix. The top left entry $x$ of $\operatorname{Sink}(A)$ satisfies $d_{6}x^{6}+d_{5}x^{5}+d_{4}x^{4}+d_{3}x^{3}+d_{2}x^{2}+d_{1}x+d_{0}=0$, where
@@
For the particular matrix $A$ in Example 1, Theorem 6 gives $091064736x^{6}-53554284x^{5}-205179237x^{4}-30568428x^{3}+51247971x^{2}+3486078x-2991816=$ 0 for the top left entry. This equation can be computed quickly with SinkhornPolynomials [18]. Dividing this equation by $243$ produces Equation (2).
@@
Let $A$ be a positive $3\times 3$ matrix, and let $a_{11}$ be the $(1,1)$ entry of $A$. If one of the $2\times 2$ minors involving $a_{11}$ is 00 and all minors not involving $a_{11}$ are not 00, then the top left entry of $\operatorname{Sink}(A)$ has degree at most $5$.
@@
Let $A$ be a positive $3\times 3$ matrix, and let $a_{ij}$ be the $(i,j)$ entry of $A$. If rows $2$ and $3$ are scalar multiples of each other, then the top left entry $x$ of $\operatorname{Sink}(A)$ satisfies $e_{3}x^{3}+e_{2}x^{2}+e_{1}x+e_{0}=0$, where

### v2

@@ Abstract.
We use a variety of computational tools to obtain a degree-$\binom{m+n-2}{m-1}$ polynomial equation conjecturally satisfied by the top-left entry of the Sinkhorn limit of a positive $m\times n$ matrix. The degree of this equation has a combinatorial interpretation as the number of minors of an $(m-1)\times(n-1)$ matrix, and the coefficients involve a determinant formula that reflects new combinatorial structure on sets of minor specifications. The tools we use include Gröbner bases, which produce equations for small matrices; the PSLQ algorithm, which produces equations for larger matrices as part of an interpolation effort that required $1.5$ years of CPU time; and ChatGPT o3-mini-high, which identified the signs of the off-diagonal entries in the determinant formula.
@@
In a 1964 paper, Sinkhorn [20] considered the following iterative scaling process. Let $A$ be a square matrix with positive entries. Scale the rows so that each row sum is $1$. Then scale the columns so that each column sum is $1$; generically, this changes the row sums. To restore the row sums $1$, scale the rows again, then scale the columns, and so on. Sinkhorn showed that the sequence of matrices obtained through this process converges to a matrix whose row and column sums are $1$ (in other words, a doubly stochastic matrix). We call this matrix the Sinkhorn limit of $A$ and denote it $\operatorname{Sink}(A)$. Sinkhorn also showed that $\operatorname{Sink}(A)$ is the unique doubly stochastic matrix $S$ with the same dimensions as $A$ such that $S=D_{1}AD_{2}$ for some diagonal matrices $D_{1}$ and $D_{2}$ with positive diagonal entries. Here $D_{1}$ can be taken to be the product of the row-scaling matrices from the iterative scaling process and $D_{2}$ the product of the column-scaling matrices.
The literature on questions related to the iterative scaling process is large; Idel’s extensive survey [10] covers results as of 2016. Mathematical applications include preconditioning linear systems to improve numerical stability, approximating the permanent of a matrix, and determining whether a graph has a perfect matching. In a number of other areas, iterative scaling was discovered independently [14, 6, 2], and it is used in machine learning to efficiently compute optimal transport distances [5]. As a result of its ubiquity and importance, many authors have been interested in fast algorithms for approximating Sinkhorn limits numerically [9, 11, 12, 15, 13, 1, 4].
@@
In particular, the top-left entry $x$ of $\operatorname{Sink}(A)$ satisfies
@@
For a general $3\times 3$ matrix $A$, Chen and Varghese [3] used numeric experiments to conjecture that the entries of $\operatorname{Sink}(A)$ generically have degree $6$ over the field generated by the entries of $A$.
@@
Let $x$ be the top-left entry of $\operatorname{Sink}(A)$. The PSLQ integer relation algorithm [8] can be used to recognize an algebraic number, given a sufficiently high-precision approximation. For the approximation
@@
In Section 3, we use the form of the equation for $3\times 3$ matrices to infer the form of the equation for $n\times n$ matrices $A$. In particular, we conjecture that the entries of $\operatorname{Sink}(A)$ generically have degree $\binom{2n-2}{n-1}$. To obtain the equation for a $4\times 4$ matrix, the Gröbner basis computation is infeasible, so instead we use PSLQ to recognize entries of Sinkhorn limits for enough integer matrices to solve for each coefficient in the equation. We also obtain several coefficients in the equation for $5\times 5$ matrices with this method. However, this does not provide enough data to identify the coefficients for $n\times n$ matrices in general.
@@
Therefore, in Section 4, we generalize to rectangular matrices by defining the Sinkhorn limit of an $m\times n$ matrix $A$ to be the matrix obtained by iteratively scaling so that each row sum is $1$ and each column sum is $\frac{m}{n}$. Again we use PSLQ, and we solve for coefficients in equations for matrices of various dimensions. We then interpolate formulas for these coefficients as functions of $m$ and $n$. By identifying combinatorial structure in these formulas, we obtain Conjecture 2 below. To state it, we introduce notation for the relevant products of minors; see Section 2 for examples.
Notation.
Define
The set $D(m,n)$ consists of the specifications of all minors of an $m\times n$ matrix $A$ that do not involve the first row or first column. We have $\lvert{D(m,n)}\rvert=\sum_{k=0}^{\min(m,n)-1}\binom{m-1}{k}\binom{n-1}{k}=\binom{m+n-2}{m-1}$. Define $A_{R,C}$ to be the submatrix of $A$ obtained by extracting the rows indexed by $R$ and the columns indexed by $C$. For each $(R,C)\in D(m,n)$, define
The minor $\Delta\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)$ involves the top-left entry $a_{11}$, and $\Gamma\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)$ is the product of $a_{11}$ and a minor that does not involve the first row or first column. This notation does not reflect the dependence on $A$, but the matrix will be clear from context. Each subset $S\subseteq D(m,n)$ specifies a monomial in the expressions $\Delta\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)$ and $\Gamma\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)$, namely
Each element of $D(m,n)$ contributes to the monomial $M(S)$ as the argument of either $\Delta$ or $\Gamma$; indeed we specify each monomial by the elements that appear as arguments of $\Delta$.
The coefficients of these monomials in Conjecture 2 are given by a determinant formula involving an adjacency-like (or possibly Laplacian-like) matrix. We define this matrix next, and we discuss the combinatorial structure behind it in Section 4.
Notation.
If $X$ is a set of positive integers and $s\in X$, define $\operatorname{pos}_{s}(X)$ to be the position, indexing from $1$, in which $s$ appears when the elements of $X$ are listed in order. For example, $\operatorname{pos}_{5}(\{2,4,5\})=3$.
Notation.
For each subset $S=\{(R_{1},C_{1}),(R_{2},C_{2}),\dots,(R_{k},C_{k})\}\subseteq D(m,n)$, let $\operatorname{adj}_{S}(m,n)$ be the $k\times k$ matrix defined as follows. The $(i,i)$ diagonal entry of $\operatorname{adj}_{S}(m,n)$ is $\lvert{R_{i}}\rvert(m+n)-mn$. For $i\neq j$, define the $4$-tuple
$\tau_{ij}=\left(R_{i}\setminus R_{j},\quad R_{j}\setminus R_{i},\quad C_{i}\setminus C_{j},\quad C_{j}\setminus C_{i}\right).$
The $(i,j)$ off-diagonal entry of $\operatorname{adj}_{S}(m,n)$ is
$\begin{cases}(-1)^{\operatorname{pos}_{s}(R_{j})+\operatorname{pos}_{t}(C_{j})}m&\text{if $\tau_{ij}=(\{\},\{s\},\{\},\{t\})$ for some $s,t$}\\ (-1)^{\operatorname{pos}_{s}(R_{i})+\operatorname{pos}_{t}(C_{i})+1}n&\text{if$\tau_{ij}=(\{s\},\{\},\{t\},\{\})$ for some $s,t$}\\ (-1)^{\operatorname{pos}_{s}(C_{i})+\operatorname{pos}_{t}(C_{j})}m&\text{if $\tau_{ij}=(\{\},\{\},\{s\},\{t\})$ for some $s,t$}\\ (-1)^{\operatorname{pos}_{s}(R_{i})+\operatorname{pos}_{t}(R_{j})}n&\text{if $\tau_{ij}=(\{s\},\{t\},\{\},\{\})$ for some $s,t$}\\ 0&\text{otherwise.}\end{cases}$
In defining $\operatorname{adj}_{S}(m,n)$, we have chosen an order on the elements of $S$. However, $\det\operatorname{adj}_{S}(m,n)$ does not depend on this order, since swapping $(R_{i},C_{i})$ and $(R_{j},C_{j})$ in the order on $S$ simply has the effect of swapping rows $i$ and $j$ and columns $i$ and $j$ in $\operatorname{adj}_{S}(m,n)$. Also note that, if we fix an order on $D(m,n)$ and use that order for each $S\subseteq D(m,n)$, then each $\operatorname{adj}_{S}(m,n)$ is a submatrix of $\operatorname{adj}_{D(m,n)}(m,n)$; this is computationally useful since it is faster to compute $\operatorname{adj}_{D(m,n)}(m,n)$ once and extract submatrices than to compute all $\operatorname{adj}_{S}(m,n)$ from the definition.
@@
Let $m\geq 1$ and $n\geq 1$. For every $m\times n$ matrix $A$ with positive entries, the top-left entry $x$ of $\operatorname{Sink}(A)$ satisfies
$\sum_{S\subseteq D(m,n)}\det\!\left(\tfrac{1}{m}\operatorname{adj}_{S}(m,n)\right)M(S)x^{\lvert{S}\rvert}=0.$
@@
The factor $\frac{1}{m}$ in Conjecture 2 arises from the asymmetry between rows and columns in our definition of $\operatorname{Sink}(A)$.
We conclude in Section 5 with several open questions. In particular, we mention that a more general iterative scaling process was introduced in 1937 by Kruithof in the context of predicting telephone traffic [14, Appendix 3d]. We conjecture that the entries of this more general limit also have degree at most $\binom{m+n-2}{m-1}$.
Our Mathematica package SinkhornPolynomials [19] uses the results of this paper to compute Sinkhorn limits rigorously for $3\times 3$ matrices and conjecturally for larger matrices.
@@ 2. The Sinkhorn limit of a $3\times 3$ matrix
It suffices to describe the top-left entry of $\operatorname{Sink}(A)$. This is because the iterative scaling process isn’t sensitive to the order of the rows or the order of the columns. If $P_{1}$ is the permutation matrix swapping rows $1$ and $i$ and $P_{2}$ is the permutation matrix swapping columns $1$ and $j$, then $P_{1}\operatorname{Sink}(A)P_{2}=\operatorname{Sink}(P_{1}AP_{2})$. In particular, the $(i,j)$ entry of $\operatorname{Sink}(A)$ is equal to the $(1,1)$ entry of $\operatorname{Sink}(P_{1}AP_{2})$.
To compute a polynomial equation satisfied by the top-left entry of $\operatorname{Sink}(A)$, we set up three matrices
@@
Let $A$ be a positive $3\times 3$ matrix. The top-left entry $x$ of $\operatorname{Sink}(A)$ satisfies $b_{6}x^{6}+\dots+b_{1}x+b_{0}=0$, where the coefficients $b_{k}$ appear in Table 1 in factored form.
@@
Table 1. Coefficients in the polynomial equation satisfied by the top-left entry of the Sinkhorn limit of a positive $3\times 3$ matrix.
In particular, the degree of each entry of $\operatorname{Sink}(A)$ is at most $6$. For the matrix $A$ in Example 1, Theorem 3 states that the top-left entry $x$ of $\operatorname{Sink}(A)$ satisfies
@@
This agrees with Equation (2). This polynomial is irreducible. Along with Theorem 3, this confirms the conjecture of Chen and Varghese [3] that the entries of $\operatorname{Sink}(A)$ for positive $3\times 3$ matrices $A$ generically have degree $6$.
Let $f(x)=b_{6}x^{6}+b_{5}x^{5}+\dots+b_{1}x+b_{0}$ be the polynomial in Theorem 3. When we project $A$ to a symmetric matrix by setting $a_{21}=a_{12}$, $a_{31}=a_{13}$, and $a_{32}=a_{23}$, then the projection of $f(x)$ factors as $-\left((a_{11}a_{23}-a_{12}a_{13})x-a_{11}a_{23}\right)^{2}g(x)$, where $g(x)$ is the degree-$4$ polynomial computed by Ekhad and Zeilberger [7] for the top-left entry.
An obvious question is whether there is a better way to write the polynomial $f(x)$ in Theorem 3. In fact there is, using determinants. We first observe that $b_{k}$ contains the factor $a_{11}^{5-k}$ for each $k\in\{0,1,\dots,5\}$. This suggests that the scaled polynomial $a_{11}f(x)$ is more natural than $f(x)$, since the coefficient of $x^{k}$ in $a_{11}f(x)$ contains the factor $a_{11}^{6-k}$ not just for $k\in\{0,1,\dots,5\}$ but for all $k\in\{0,1,\dots,6\}$.
The leading coefficient $a_{11}b_{6}$ of $a_{11}f(x)$ is the product of $6$ minors of $A$. More specifically, it is the product of all the minors of $A$ that involve $a_{11}$. Furthermore, the constant coefficient $a_{11}b_{0}$ of $a_{11}f(x)$ is the product of $a_{11}^{6}$ and the $6$ minors of $A$ that do not involve the first row or first column (including the determinant $1$ of the $0\times 0$ matrix).
@@
We have seen that this is the case for $a_{11}b_{6}$ and $a_{11}b_{0}$. Remarkably, it turns out that $a_{11}b_{5},a_{11}b_{4},\dots,a_{11}b_{1}$ can be written this way as well. We use the notation $\Delta$, $\Gamma$, and $M$ established in Section 1. To make a subset $S=\{(R_{1},C_{1}),(R_{2},C_{2}),\dots,(R_{k},C_{k})\}$ of $D(m,n)$ easier to read, we format it as
@@
We continue to let $A$ be the general $3\times 3$ matrix in Equation (3), and let $x$ be the top-left entry of $\operatorname{Sink}(A)$. To rewrite the coefficients $a_{11}b_{k}$, we look for coefficients $c_{S}\in\mathbb{Q}$, indexed by subsets $S\subseteq D(3,3)$, such that
@@
Therefore the expression for $b_{3}$ in Table 1 does not uniquely determine the coefficients $c_{S}$ where $\lvert{S}\rvert=3$. However, there is additional information we can use to obtain uniqueness. The polynomial $a_{11}f(x)$ possesses symmetries arising from the following invariance properties.
Since the iterative scaling process isn’t sensitive to row order, the top-left entry of $\operatorname{Sink}(A)$ is invariant under row permutations of $A$ that fix the first row. Similarly for column permutations. Additionally, the top-left entry of $\operatorname{Sink}(A)$ is invariant under transposition of $A$; this follows from Sinkhorn’s result that $\operatorname{Sink}(A)$ is the unique doubly stochastic matrix $S$ such that $S=D_{1}AD_{2}$ for some diagonal matrices $D_{1}$ and $D_{2}$. This suggests the following equivalence relation.
@@
Motivated by Example 5, we add the constraint that $c_{T}=c_{S}$ when $T\equiv S$, so that the coefficients $c_{S}$ reflect the symmetries of $a_{11}f(x)$. With this constraint, the coefficient $a_{11}b_{3}$ has a unique representation as a linear combination of monomials $M(S)$ where $\lvert{S}\rvert=3$. Writing each coefficient in Table 1 using class sums gives the following improvement of Theorem 3. Here $d_{k}=a_{11}b_{k}$.
@@
Let $A$ be a positive $3\times 3$ matrix. The top-left entry $x$ of $\operatorname{Sink}(A)$ satisfies $d_{6}x^{6}+d_{5}x^{5}+d_{4}x^{4}+d_{3}x^{3}+d_{2}x^{2}+d_{1}x+d_{0}=0$, where
@@
For the particular matrix $A$ in Example 1, Theorem 6 gives $091064736x^{6}-53554284x^{5}-205179237x^{4}-30568428x^{3}+51247971x^{2}+3486078x-2991816=$ 0 for the top-left entry. This equation can be computed quickly with SinkhornPolynomials [19]. Dividing this equation by $243$ produces Equation (2).
@@
Let $A$ be a positive $3\times 3$ matrix, and let $a_{11}$ be the top-left entry of $A$. If one of the $2\times 2$ minors involving $a_{11}$ is 00 and all minors not involving $a_{11}$ are not 00, then the top-left entry of $\operatorname{Sink}(A)$ has degree at most $5$.
@@
Let $A$ be a positive $3\times 3$ matrix, and let $a_{ij}$ be the $(i,j)$ entry of $A$. If rows $2$ and $3$ are scalar multiples of each other, then the top-left entry $x$ of $\operatorname{Sink}(A)$ satisfies $e_{3}x^{3}+e_{2}x^{2}+e_{1}x+e_{0}=0$, where